What is the molality of a solution made by dissolving 50 g of acetic acid

(CH3COOH) in 950 g of ethanol.


1
Expert's answer
2021-03-03T01:59:53-0500

Ans:- Molecular mass of "CH_3COOH" is "60" "\dfrac{g}{mol}"

Now find the moles of acetic acid

"\Rightarrow" moles of acetic acid"=\dfrac{given mass}{molecular mass}"

"\Rightarrow" "\dfrac{50}{60}=0.833mol"


Therefore Molality"=\dfrac{no. of moles of solute}{mass of solvent in (gm)}\times1000"

"\Rightarrow" molality of acetic acid"=\dfrac{0.833}{950}\times1000=0.8771m"


Therefore molality of acetic acid in ethanol is "0.8771m"